a beam of light travelling in air strikes a glass table top at an angle of 44.5 degrees and the glass has an index of refraction of 1.5.
what us the angle of refraction for the light transmitted into the glass
1
Expert's answer
2019-11-05T10:59:24-0500

According to Snellus" law, we can write

"\frac {\sin{44.5°}}{\sin{α}}=1.5 (1)"

Using (1) we get

"\sin{α}=0.4673"

α=28°


Answer

28°